OpenStudy (anonymous):

What is the approximate perimeter of a semicircle with a radius of 8 m? Use pi star times = 3.14. Express your final answer to the nearest hundredth. PLEASE HELP! DON'T GUESS! I WILL GIVE MADAL!

OpenStudy (anonymous):

@Bookworm14

OpenStudy (anonymous):

hey up123

OpenStudy (anonymous):

hey, nice pic

OpenStudy (anonymous):

@Bookworm14

OpenStudy (anonymous):

thanks found it in library dad did it thought cool

OpenStudy (anonymous):

nice, Hi Book worm 14!

OpenStudy (anonymous):

could you please help me?

OpenStudy (anonymous):

so first find the circumference 2 times pi and radius

OpenStudy (bookworm14):

Hi! and yes I will :)

OpenStudy (bookworm14):

what do you believe it is?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

25.12

OpenStudy (anonymous):

so 2*3.14*8 basically because in order to find the circumference you need to find 2 pi radius

OpenStudy (anonymous):

is my answer, i want to be sure

OpenStudy (anonymous):

and then divide by two to get 25.04

OpenStudy (anonymous):

that what i got

OpenStudy (anonymous):

k12.http.internapcdn.net/k12_vitalstream_com/CURRICULUM/329902/CURRENT_RELEASE/MS_PA_S2_03_09_WHB_circumference.htm

OpenStudy (anonymous):

opps, not a link

OpenStudy (bookworm14):

so for this you will find you circumference which is 2 *8 * pi this equals 50.24 a semi-circle is half of a circle so 50.24/ 2= 25.12 Your Correct! :D

OpenStudy (anonymous):

when I divided, i got 12.56

OpenStudy (anonymous):

weird

OpenStudy (anonymous):

really, YAY

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

i got 25.04 weird

OpenStudy (anonymous):

use google calculator, it helps :)

OpenStudy (bookworm14):

Your welcome and @ekim5095 did you use the pi button on a calculator?

OpenStudy (anonymous):

wow really i used paper

OpenStudy (anonymous):

lol 12 year old using aper old fashioned yeah

OpenStudy (anonymous):

paper

OpenStudy (bookworm14):

I am 15 and I hate calculators so I almost always use pencil and paper unless i have to use the calc :)

OpenStudy (anonymous):

lol same

OpenStudy (anonymous):

I am 14, not to far off from u ;)

OpenStudy (anonymous):

same, sometimes, but same

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

perimeter of a semi circle=pi*radius+Diameter =3.14*8+16 =25.12+16 =41.12 meter.

OpenStudy (anonymous):

^^^ He's right I just took the test for anyone looking for answers